I have the answer to the following problem in my book, but I can't figure out the formula to use to get it right. Please help me figure out the correct formula. Thanks so much.:
An airplane flying against the wind travels 300 miles in the same time that it would take the same plane, flying the same speed, to travel 400 miles with the wind. If the wind speed is 20 mph, find the speed of the airplane in still air.
I tried this: 300/N-20 = x, 400/N+20 = x But I could not solve for x with just a number because of the N. This was Distance / Rate = Time. I also tried: 300 = (x)(N-20), 400 = (x)(N+20), but again cannot solve completely with two unknowns. Please help!!!!

Answer by scott8148(6628)

you don't need x __ the problem states "in the same time"

so 300/N-20=400/N+20

300N+6000=400N-8000 __ 14000=100N